OpenStudy (anonymous):

15b^2 - 19b -10 factor by grouping

OpenStudy (anonymous):

(3b-5) (5b+2)

myininaya (myininaya):

we have Ab^2+Bb+C find two factors of A*C that multiply to be A*C and add up to be B A*C=15*(-10)=5(-30)=75(-2)=25(-6)=-25(6) B=-19=-25+6 so replace Bb=-19b with -25b+6b so we have 15b^2-25b+6b-10 now factor by grouping 5b(3b-5)+2(3b-5) (3b-5)(5b+2)

OpenStudy (anonymous):

thank you guys! sorry it took so long to respond.